# Improving Code Reviewer Recommendation: Accuracy, Latency, Workload, and Bystanders

### v1

@@ 6. Balancing Reviewer Workload
Condition Workload Metric Top1 Accuracy Top1 Workload Top3 Accuracy Top3 Workload
@@
Production Time Spent -4.80 -18.06 -5.34 -2.84
Engineers have different skillsets which will will always skew reviewer workload (Rigby and Storey, 2011), it is still interesting to understand if a recommender that is aware of workload might distribute the load of reviewing and potentially drive down cycle time by suggesting the least busy expert. For example, Microsoft weighted their code reviewer recommender by a candidate’s workload and reduced the time it takes for a review (Asthana et al., 2019). A similar experiment at Ericsson (Strand et al., 2020) balanced reviewer workload, but was unsuccessful and was rolled back. As a result, we build upon the Microsoft work and we take the score from RevRecV2 that is now in production and weight it by the workload of a candidate as described in Section 4.2.
We experiment with three types of workload weights the number of upcoming meetings in the next 7 days, the reviewer activity in the last 7 days (Asthana et al., 2019), and time spent reviewing in the last 7 days. We calculate each over a 7 day period because developer patterns on a particular day tend to vary, but the entire work week tends to be more predictable (Chen et al., 2022). We conduct a historical backtest to determine which measure works as well as a workload weight. We then roll out a workload in production and determine if it is effective.
Our goal metric is to reduce the median amount of time spent reviewing of the reviewer who made an action on a diff. We expect a decrease in accuracy, however, because we only re-rank the top 5 reviewers, the same reviewers will be shown they will just be ordered by their workload, so top5 precision will be unchanged.
Our other guardrail metrics are the same is in the previous section. We do not not want to increase review cycle time, we do more quickly reviewed diffs, so the time spent per diff should remain unchanged, we do not want to see fewer overall clicks on our recommendations, and we do not want latency to increase.
6.1. Historical Backtest
Most of the code review recommender literature uses historical data to simulate how well the recommender would have done at recommending the person who actually does the review (Thongtanunam et al., 2015; Mirsaeedi and Rigby, 2020). This accuracy is often measured with TopN accuracy. We also measure the median workload of the top recommended candidate. The top portion of Table 6, shows the results from 1.2 million historical diff reviews that were performed early 2023. We see that the Top1 accuracy decreases dramatically using workload balancing by around -20% compared to RevRecV2. The workload differences are less consistent, with upcoming meetings performing poorly, and review activity and time spent performing much better with respective decreases of -57% and -73%. The Top3 accuracy shows similar patterns with time spent having the largest drop in workload.
6.2. RevRecWL in Production
Since we are only re-ordering the reviewers suggested by RevRecV2 and are still displaying the Top5 recommended reviewers, we decided to use the most aggressive weight metric, TimeSpent7d, because it had the largest drop in workload. The unit in the experiment was the author of a diff, so each other would either have RevRecV2 or RevRecWL. We had 14k diff authors in each condition for a total of 28k authors.
@@ 6.2. RevRecWL in Production
We were surprised how different the results from the backtest were from the actual people who performed the review, see the last row in Table 6. When the author selected the top recommendation, the accuracy decreased much less than in the back test -5% vs -21% decrease in accuracy and the workload only decreased -18% vs -73% in the production vs backtest. A similar discrepancy can be seen with Top3, recall that Top5 will be unchanged.
@@
Our guardrail metric, number of clicks explains the discrepancy, we saw a decrease in selection of the Top1 candidate by 30% of all diffs to 20%. Overall, however, the difference in the number of clicks did not change in a statistically significant manner. What this means is that when someone with low workload was an inappropriate reviewer, the author simply ignored them and moved lower in the list of recommended reviewers. This results also reinforces how important it is to first model expert reviewers with RevRecV2 and then simply re-rank them.
@@
The remaining guardrail metrics latency, time spent per diff, and review cycle time where not changed in a statistically significant manner. We were unable to replicate the Microsoft (Asthana et al., 2019) improvement in review interval.
@@
When a reasonable candidate with lower workload was available we saw a large reduction in workload. When an inappropriate candidate was recommended, simply because they had low workload, the author looked down the list for a more experienced developer, and we saw a drop in the top ranked candidate being selected. We did not see any statistically significant change in review cycle time or latency. As a result, we decided to roll RevRecWL out to 100% at Meta.
@@ 7. The Bystander Effect
When a task is assigned to a group of people, the likelihood of any individual acting is lower than when an individual is explicitly assigned to this task (Fischer et al., 2011). At Meta, diff authors can assign teams and groups of developers rather than selecting individual reviewers. In some cases, an author might not mind which developer on a team reviews a diff. Furthermore, Meta uses a monorepo where developers often make changes that crosscut the codebase and are owned by another development team. The author might not know who the best reviewer is and will instead select a team. Selecting a reviewer group creates the risk of each member thinking, “someone else will probably review this.” We decided to leverage the wins from our reviewer recommendation investments last year to assign a single individual reviewer to diffs where only reviewer groups were assigned in an attempt to reduce the Bystander Effect.
We developed and tested two strategies to explicitly assign a reviewer on diffs where the author did not choose an individual reviewer. (1) BystanderRecWL, automatically assigns the top candidate reviewer from RevRecWL to the diff, i.e. the top expert with the current lowest workload. (2) BystanderRecRnd, the simpler of the two, built upon RevRecV2 and simply chose one of the top three reviewers at random. Our goal is that reviewers will feel a sense of ownership for diffs where they were assigned as a reviewer, and review it more quickly, thus reducing the time-in-review metric. We do not want to see regressions in how long a developer views a diff. We do not track latency, clicks, or workload as they are unlikely to be impacted by this experiment.
7.1. Results for Bystander Effect
We had expected BystanderRecWL to outperform the simpler
BystanderRecRnd because we figured that a low workload person would respond more quickly. We were surprised to find that BystanderRecWL had no statistically significant impact on any of the metrics including our goal metric of reducing time-in-review.
In contrast, BystanderRecRnd, was extremely successful and reduced time-in-review by -11.6%. There was no statistically significant change in the guardrail metric of the amount of time reviewers spent on each diff. We rolled BystanderRecRnd out to 100% of diffs at Meta.
@@ 7.1. Results for Bystander Effect
The bystander effect occurs in code review when only a team is assigned to a review a diff instead of an individual. At Meta we randomly assigned one of the top3 experts to review the diff and found a drop of -11.6% in time-in-review with no regressions in our guardrail metrics. We rolled BystanderRecRnd out to 100% of engineers.
@@ 8.1. Generalizability
Drawing general conclusions from empirical studies in software engineering is difficult because any process depends on a potentially large number of relevant context variables. The analyses in the present paper were performed at Meta, and it is possible that results might not hold true elsewhere. We cannot release our data, even in an anonymized format, because it would violate legally binding employee privacy agreements. However, our study does cover a very wide swath of software engineering. The software systems covers millions of lines of code and 10’s of thousands of developers who are both collocated and working at multiple locations across the world. We also cover a wide range of domains from user facing social network products and virtual and augmented reality projects to software engineering infrastructure, such as calendar, task, and release engineering tooling.
We have discussed the many reviewer recommenders available in the literature above. Instead of reimplementing these recommenders, we have used the feature categories suggested by Al-Zubaidi et al. (Al-Zubaidi et al., 2020), and included features from each categories that would not increase latency in our model. It is possible that smaller organizations or projects can use code review social network features etc, unfortunately they do not work at our scale.
@@ 8.2. Construct Validity
We have also reviewed the literature to ensure that we covered the normal features/constructs include in reviewer recommenders. We have used those that can be extracted from Meta’s history and those that can be calculated in a timely manner.
The implicit assumption when using historical data is that the person who performs an action on a diff is a good reviewer. Recommenders have been trained on this dataset (Kovalenko et al., 2018). However, as we saw in, for example, our RevRecWL evaluation, when recommenders that have alternative purposes, i.e. spreading workload, the results can be drastically different in a historical backtest compared to an actual release in production. To ensure that we cover a wide range of constructs, we use multiple goal and guardrail measures in each experiment and do not just rely on results from a backtest.
@@ 8.3. Internal Validity
Each of the recommenders presented in this paper have been tested in a production experiment based on goal and guardrail metrics. However, each experiment was run at a different time, with RevRecV2 experimentation in the Spring of 2022, RevRecWL in winter of 2023, and BystanderRecRnd in Spring of 2023. We cannot report the exact time-frames, but report the number of diffs or diff authors so that the size of sample can be evaluated by the reader. As part of engineering at Meta we track metrics across time and did not see any changes in accuracy, time-in-review, and TimeSpent outside the experiments that we ran.
@@ 9. Literature and Discussion
The first reviewer recommendation system we are aware of was introduced by Balachandran et al. (Balachandran, 2013). They used authorship of the changed lines in a code review (using git blame) to identify who had worked on that code before and suggested a ranked list of this set as potential reviewers. Similarly, Thongtanunam et al. (Thongtanunam et al., 2015) proposed RevFinder, a reviewer recommender based on file locations. RevFinder is able to recommend reviewers for new files based on reviews of files that have similar paths in the filesystem. The approach was evaluated on over 40,000 code reviews across three OSS projects, and recalls a correct reviewer in the top 10 recommendations 79% of the time on average. Like RevRecV1 these approaches would suggest reviewers who had moved onto other projects because they had edited the files in the past.
There has been a lot of interest in consuming features beyond file history when building reviewer recommendation models. Rahman et al. (Rahman et al., 2016) propose CORRECT an approach to recommend reviewers based on their history across all of GitHub as well as their experience certain specialized technologies associated with a pull request. Jiang et al. (Jiang et al., 2017) examine the impact of various attributes of a pull request on a reviewer recommender, including file similarity, PR text similarity, social relations, and “activeness” and time difference. They find that adding measures of activeness to prior models increases performance considerably. Zhang et al. (Zhang et al., 2022) built a Socio-Technical graph to capture the social interactions among authors and reviewers. Further, they use Graph Convolution Networks (GCNs) to encode the social interaction information into high-dimensional space, and formulate reviewer recommendation as a link prediction problem. We describe the features that we were able to include in our model in Table 2 building upon the categories suggested by Al-Zubaidi et al. (Al-Zubaidi et al., 2020).
Latency and Performance. Some of these recommenders use social network features and other features that required substantial processing. As noted by Microsoft (Asthana et al., 2019; Zanjani et al., 2016) and Ericsson (Strand et al., 2020), these cannot be used in production systems due to processing delay. However, they did not report results on the importance of latency. Interestingly, when we reduced the latency from RevRecV1 to RevRecV2 we saw that lower latency resulted in authors seeing recommendations earlier before they had decided on who would do the review and many more clicks on recommendations. Future works should provide information on prediction latency.
@@ 9.1. Reviewer Workload and Recommenders
Kovalenko et al. (Kovalenko et al., 2018) found that expertise based recommenders tend to suggest reviewers that are always obvious to authors 52% and 63% of the time at Microsoft and JetBrains respectively. They are usually obvious to authors 37% and 38% of the time. As a result, recommendations often clicked on because it is easier to click than type a reviewer’s name. Another interesting finding was that authors often found that outdated recommendations were made based on people who use to work on the files. Authors wanted more information about the current knowledge level of the candidate reviewers as well as how busy, i.e. what the workload of the reviewers are.
@@
Al-Zubaidi et al. (Al-Zubaidi et al., 2020) treat reviewer recommendation as a multi-objective search problem where they maximize for the chance of participating in review and minimze the shannon entropy, i.e. skewness, of the people selected. Their recommender considers review workload, but has low precision ranging from 16% to 20% depending on the open source project. They do not use a measure of workload as an outcome, so it is unclear if their approach succeeds in spreading workload. They do not report on latency or performance, and it is unclear if this search-based approach could be used at large scale companies.
Strand  et al. (Strand et al., 2020) evaluate a workload aware recommender at Ericsson. In a backtest they found that a recommender aware of the file change history and the level of recent workload activity was able to make recommendations with an MRR of .37 or an average rank of 2.7. Despite these reasonable recommendations, after manually evaluating the performance of the recommender on 47 changes they determined that it did not reduce lead time or workload for code review and they decided to not rollout the recommender.
Murphy-Hill et al. (Murphy-Hill et al., 2023), examine workload at Google and discover an systemic imbalance in gender and code review selection. Authors manually select men more often than women as reviewers. Since recommenders are trained on prior data, the recommenders contain this bias. In our work, we do not explicitly factor gender into our recommender, however, we re-rank reviewers based on workload, so if women are assigned fewer reviewers, they will automatically be assigned more reviews as workload is balanced. We also use random in our bystander experiment, which Murphy-Hill et al. finds balances the gender.
10. Concluding Remarks
We make the following contributions related to accuracy and latency, workload, and the bystander effect:

### v2

@@ 6.1. Historical Backtest Results
Method Workload Metric Top1 Accuracy Top1 Workload Top3 Accuracy Top3 Workload
@@
Expt. 2 Time Spent -4.80 -18.06 -5.34 -2.84
6.2. Results for Expt 2. RevRecWL in Production
We discussed our experimentation framework and metrics in background Section 4.1. To summarize, our goal metric is to reduce the workload of the Top recommended reviewers. Our backtest indicates that we will see a drop in the accuracy. This reduction is the tradeoff for a more evenly distributed workload. Specifically, we expect the accuracy to drop by no more than -20 percentage points. Our other guardrail metrics remain the same: we do not want to see fewer clicks on our recommendations, we do not want people to spend less time reviewing the diff, and we do not expect to see any change in how long reviews are waiting to be reviewed. The unit in the experiment was the author of a diff, so each author would either have RevRecV2 or RevRecWL. We had 14k authors in each condition for a total of 28k authors.
Table 9. For Expt. 2 on workload balancing, we had RevRecV2 as the control and RevRecWL as the test. The unit was 28k diff authors. We conducted the experiment in Winter 2023. From our backtest in Table 8, we expected to see a drop in reviewer workload as reviews become more evenly distributed. From the backtest shown in Table 8, we also expected to see a trade-off/drop in accuracy of up to $-21$ pp. The table shows that we did not see as much of a drop in workload, but we also did not see the same drop in accuracy. We did see an unexpected drop in Top 1 Clicks, however, the overall clicks remained constant.
Outcome Metric Type Expectation Change p-value
Workload Top 1 Goal Decrease $-18.06$ pp. $p\ll 0.01$
Workload Top 3 Goal Decrease $p=0.19$
Accuracy Top 1 Guardrail Decrease must be $>-21$ pp. $-4.90$ pp. $p\ll 0.01$
Accuracy Top 3 Guardrail Decrease must be $>-20$ pp. $-5.34$ pp. $p\ll 0.01$
Clicks Top 1 Guardrail No Decrease $-10$ pp. $p\ll 0.01$
All Clicks Guardrail No Decrease $p=0.09$
TimeInReview Guardrail No Increase $p=0.75$
TimeSpent Guardrail No Decrease $p=0.85$
@@
Results: We were surprised how different the results from the backtest were from the controlled production experiment, see the last row in Table 8. When the author selected the top recommendation, the accuracy decreased much less than in the backtest -5% vs -21% decrease in accuracy and the workload only decreased -18% vs -73% in the production vs backtest. The workload drop at Top3 was not statistically significant.
@@
Our guardrail metric, number of clicks explains the discrepancy, we saw a decrease in selection of the Top1 candidate by 30% of all diffs to 20%. Overall, however, the difference in the number of clicks did not change in a statistically significant manner. What this means is that when someone with a low workload was an inappropriate reviewer, the author simply ignored them and moved lower in the list of recommended reviewers. This result also reinforces how important it is to first model expert reviewers with RevRecV2 and then simply re-rank.
@@
The remaining guardrail metrics time spent per diff and review cycle time were not changed in a statistically significant manner. This means that we were unable to replicate the Microsoft (Asthana et al., 2019) reduction in the time pull-requests spent in review.
@@
When a reasonable candidate with lower workload was available, we saw a large reduction in workload. When an inappropriate candidate was recommended, simply because they had low workload, the author looked down the list for a more experienced developer, and we saw a drop in the top ranked candidate being selected. We did not see any statistically significant change in overall clicks or review cycle time. Since there was a reduction in workload for the Top1 reviewer, i.e. a more evenly distributed workload, we replaced RevRecV2 with RevRecWL and rolled it out to 100% of diffs at Meta.
@@ 7. The Bystander Effect
In psychology, the ”bystander effect” occurs when a task is assigned to a group of people and leads to a lower likelihood of any individual acting than when an individual is explicitly assigned to a task (Fischer et al., 2011). At Meta, diff authors can assign teams and groups of developers rather than selecting individual reviewers. In some cases, an author might not mind which developer on a team reviews a diff. Furthermore, Meta uses a monorepo where developers often make changes that crosscut the codebase and are owned by another development team. The author might not know who the best reviewer is and will instead select a team. Selecting a reviewer group creates the risk of each member thinking, “someone else will probably review this.” We decided to leverage the wins from our reviewer recommendation investments last year to assign a single individual reviewer to diffs where only reviewer groups were assigned in an attempt to reduce the Bystander Effect.
The code review team decided to take a simple randomization strategy to overcome the bystander effect. BystanderRecRnd, is built directly upon RevRecV2, and we simply chose one of the top three reviewers at random. Our goal is that reviewers will feel a sense of ownership for diffs where they were assigned as a reviewer, and review it more quickly, thus reducing the TimeInReview metric. We do not want to see regressions in how long a developer views a diff. We do not track latency, clicks, or workload as they are unlikely to be impacted by this experiment. Our experimental unit was 12.5k authors that were evenly divided between the control and test conditions.
Results. BystanderRecRnd, was extremely successful and reduced TimeInReview by -11.6% with $p\ll 0.01$. There was no statistically significant change in the guardrail metric of the amount of time reviewers spent on each diff. We rolled BystanderRecRnd out to 100% of diffs at Meta where there was only a team or group assigned the review.
Table 10. Expt. 3: the bystander effect. We compared the TeamOnly assignment control to BystanderRecRnd test. The experiment involved 12.5k authors evenly divided between the two conditions in Spring 2023. The bystander effect is real, and it was effectively minimized by randomly assigning one of the top three recommended reviewers.
Outcome Metric Type Expectation Change p-value
[…]
TimeSpent Guardrail No Decrease $p=0.37$
@@
The bystander effect occurs in code review when only a team is assigned to review a diff instead of an individual. At Meta we randomly assigned one of the top three experts to review the diff and found a drop of -11.6% in TimeInReview ($p\ll 0.01$), with no statistically significant regressions in how long reviewers spent per review, i.e. TimeSpent. We rolled BystanderRecRnd out all diffs that only had a team assigned, and they will now also have an individual assigned.
@@ 8.1. Generalizability
Drawing general conclusions from empirical studies in software engineering is difficult because any process depends on a potentially large number of relevant context variables. The analyses in the present paper were performed at Meta, and it is possible that results might not hold true elsewhere. We cannot release our data, even in an anonymized format, because it would violate legally binding employee privacy agreements. However, the software systems under study cover millions of lines of code and 10s of thousands of developers who are both collocated and working at multiple locations across the world. We also cover a wide range of domains from user facing social network products and virtual and augmented reality projects to software engineering infrastructure, such as calendar, task, and release engineering tooling.
There are many reviewer recommenders available in the literature above, and instead of reimplementing these recommenders, we have used the feature categories suggested by Al-Zubaidi et al. (Al-Zubaidi et al., 2020), and included features from each categories that would not increase latency in our model. We have used those that can be extracted from Meta’s history and those that can be calculated in a timely manner. It is possible that smaller organizations or smaller projects can use computationally expensive features, but unfortunately they do not work at our scale.
@@ 8.2. Construct Validity
The implicit assumption when, using historical data to assess TopN accuracy, is that the person who performed an action on a diff is a good reviewer. In our backtest, we saw that the TopN accuracy can be quite different from the person who actually conducts the review in a controlled experiment in production. Virtually all prior work has assumed that authors would accept the recommendations, our work sheds doubt on historical benchmarks as the gold standard of accuracy.
In our experiments, we used the outcome measures that are commonly used by the code review team at Meta. The two most commonly used measures are TimeInReview and TimeSpent. Most works capture the wall time from when the pull-request is published until it is merged or abandoned. In contrast, at Meta we only capture the time that a diff is actively being reviewed or is waiting for review measures. Clicks and latency are uncontroversial measures that are simple to calculate.
Engineers have a variety of tasks and in this work we only backtested the number of meetings, reviewing activity, and the amount of time spent viewing diffs as measures of workload. In the end we selected TimeSpent, but we look forward to seeing what type of workload is effective at other companies.
@@ 8.3. Internal Validity
Each of the recommenders presented in this paper has been tested in a production experiment based on goal and guardrail metrics. However, each experiment was run at a different time, with RevRecV2 experimentation in the Spring of 2022, RevRecWL in Winter of 2023, and BystanderRecRnd in Spring of 2023. We cannot report the exact time-frames, but report the number of diffs or diff authors so that the size of the sample can be evaluated by the reader. As part of engineering at Meta we track metrics across time and did not see any changes in accuracy, TimeInReview, and TimeSpent outside the experiments that we ran.
While the accuracy of the recommender is unaffected by collinearity among the features, the relative feature importance can be impacted. For example, we have the feature of the total amount of time a reviewer examined an author’s diff and the time they did this when they were explicitly assigned to review the author’s diffs. These features are definitely correlated and one is even a subset of the other. We kept both features because the engineering team thought that the features were both interesting, and as can be seen in the table, both add predictive power to the model. While the exact feature importance might vary depending on the project or environment, we based our feature selection on categories of features found in the literature and feel that they are a robust representation of features that are useful in reviewer recommendation.
8.4. Conclusion Validity
Conclusion Validity (Wohlin et al., 2012) is defined as the degree to which conclusions reached (e.g. about relationships between factors) are reasonable within the data collected. We use experiments that randomize treatments, thus eliminating researcher bias.
We do not explicitly control for factors including the size and complexity of the diff (and of the underlying codebase), author and reviewer expertise, programming language, type of functionality being implemented, time of the year, and day of the week. Instead we use randomization to help control for observed and unobserved variables, and we also have large sample sizes.
We report all experiments that were conducted and the results for all goal metrics (whether statistically significant or not). The p-values are adjusted for multiple comparisons in case more than one goal metric was used. Finally, since we only compare the distribution of response variables to a single predictor (experimental condition), we minimize the subjectivity due to researchers selecting one versus another subset of predictors for reporting.
@@ 9. Literature and Discussion
We interleave our discussion with the literature showing how we advance the state-of-the-art reviewer recommenders.
9.1. General Reviewer Recommenders
The first reviewer recommendation system we are aware of was introduced by Balachandran et al. (Balachandran, 2013). They used authorship of the changed lines in a code review (using blame) to identify who had worked on that code before and suggested a ranked list of this set as potential reviewers. Similarly, Thongtanunam et al. (Thongtanunam et al., 2015) proposed RevFinder, a reviewer recommender based on file locations. RevFinder is able to recommend reviewers for new files based on reviews of files that have similar paths in the filesystem. The approach was evaluated on over 40,000 code reviews across three OSS projects, and recalls a correct reviewer in the Top10 recommendations 79% of the time on average. Like RevRecV1 these approaches would suggest reviewers who had moved onto other projects because they had edited the files in the past.
There has been a lot of interest in consuming features beyond file history when building reviewer recommendation models. Rahman et al. (Rahman et al., 2016) propose CORRECT an approach to recommend reviewers based on their history across all of GitHub as well as their experience with certain specialized technologies associated with a pull request. Jiang et al. (Jiang et al., 2017) examine the impact of various attributes of a pull request on a reviewer recommender, including file similarity, PR text similarity, social relations, and “activeness” and time difference. They find that adding measures of activeness to prior models increases performance considerably. Zhang et al. (Zhang et al., 2022) built a Socio-Technical graph to capture the social interactions among authors and reviewers. Further, they use Graph Convolution Networks (GCNs) to encode the social interaction information into high-dimensional space, and formulate reviewer recommendation as a link prediction problem. We describe the features that we were able to include in our model in Table 1 building upon the categories suggested by Al-Zubaidi et al. (Al-Zubaidi et al., 2020).
9.2. Latency and Performance.
Some of these recommenders use social network features and other features that require substantial processing. As noted by Microsoft (Asthana et al., 2019; Zanjani et al., 2016) and Ericsson (Strand et al., 2020), these cannot be used in production systems due to processing delay. We are unaware of any paper that reports results on latency of reviewer recommendations. Interestingly, when we reduced the latency from RevRecV1 to RevRecV2 we saw that lower latency resulted in authors seeing recommendations earlier before they had decided on who would do the review and many more clicks on recommendations. Future works should provide information on prediction latency.
@@ 9.3. Reviewer Workload and Recommenders
Kovalenko et al. (Kovalenko et al., 2018) found that expertise based recommenders tend to suggest reviewers that are always obvious to authors 52% and 63% of the time at Microsoft and JetBrains respectively. As a result, authors often clicked on because it is easier to click than type a reviewer’s name. Another interesting finding was that authors often found that outdated recommendations were made based on people who use to work on the files. Interviewed developers wanted more information about the current knowledge level of the candidate reviewers as well as how busy, i.e. what the workload of the reviewers are.
@@
Al-Zubaidi et al. (Al-Zubaidi et al., 2020) treat reviewer recommendation as a multi-objective search problem where they maximize for the chance of participating in review and minimize the shannon entropy, i.e. skewness, of the people selected. Their recommender considers review workload, but has low precision ranging from 16% to 20% depending on the open source project. They do not use a measure of workload as an outcome, so it is unclear if their approach succeeds in spreading workload. They do not report on latency or performance, and it is unclear if this search-based approach could be used at large scale companies.
Strand  et al. (Strand et al., 2020) evaluate a workload aware recommender at Ericsson. In a backtest they found that a recommender aware of the file change history and the level of recent workload activity was able to make recommendations with an MRR of 0.37 or an average rank of 2.7. Despite these reasonable recommendations, after manually evaluating the performance of the recommender on 47 changes they determined that it did not reduce lead time or workload for code review and they decided to not roll-out the recommender. Like Ericsson, we did not find a change in cycle time, however, we did find changes in the workload of the reviewer.
Hajari  et al. (Hajari et al., 2024) used the Gini coefficient to capture how unequal workload is and on open source projects found a near perfect Pareto principle with the top 20% of reviewers doing 80% of the reviews. The authors simulate the change in Gini over multiple years of data and conclude that a workload aware recommender similar to those used by Microsoft (Asthana et al., 2019) can reduce workload. Our work threatens the conclusions of this paper because workload balanced recommendations can be ignored with the “normal” experts still doing most of the work. It is interesting future work to see if the Gini co-efficient changes over the period of multiple years when a workload balanced recommender is released in production.
Murphy-Hill et al. (Murphy-Hill et al., 2023), examine workload at Google and discover a systemic imbalance in gender and code review selection. Authors manually select men more often than women as reviewers. Since recommenders are trained on prior data, the recommenders contain this bias. In our work, we do not explicitly factor gender into our recommender, however, we re-rank reviewers based on workload, so if women are assigned fewer reviewers, they will automatically be assigned more reviews as workload is balanced. We also use random in our bystander experiment, which Murphy-Hill et al. finds balances the gender of recommended reviewers.
10. Conclusion
At Meta we make the following contributions related to accuracy and latency, workload, and the bystander effect: